# Penrose limits and TsT for fibered $I$-branes

### v2

@@ Abstract
In this paper we analyze a generalized ”single-trace $T\bar{T}$” deformation, defined by a TsT transformation, of the fibered $I$-brane solution from [1]. We use the Penrose limit to understand it, and we consider both the TsT followed by the Penrose limit, as well as the Penrose limit followed by TsT. We describe the spin chains obtained in field theory. In the first case we find that, indeed, the TsT transformation preserves solvability in a simple way, as in the standard $T\bar{T}$ case. In the second case, we have several options, but none is simple enough to be conclusive, however, one case gives us an asymptotically free and IR nontrivial field theory sector, and another a new parallelizable pp wave.
@@ 2 Review of $I$-brane background(s)
The supersymmetry of this background was studied in [1], where it was found that it preserves four supersymmetries when the parameters $e_{A,B},m,Q_{A,B}$ satisfy $m=0,\ e_{A}Q_{B}=\pm e_{B}Q_{A}$.
@@ 3.1 TsT transformation ($T\bar{T}$ deformation) in gravity and gauge theory
The general rules for T duality (both for the NSNS and the RR fields) are known from the literature, see [20], [21], [22], but for the reader’s convenience, the general rules are given in the Appendix A.2. Since the TsT transformation acts on $x$ and $t$, in the intermediate steps we rewrite only the parts of the metric affected by these transformations. Performing first a T-duality along $x$, we obtain
$ds^{2}$ $=$ $-rdt^{2}+\frac{1}{r}dx^{2}+\cdots.$ (3.1)
@@
$ds^{2}=-rdt^{2}-2\eta rdtdx+\left(\frac{1-\left(r\eta\right)^{2}}{r}\right)dx^{2}+\cdots\;,$ (3.3)
@@
Finally we T-dualize back in $x$. The NSNS part of the background is
$ds_{st}^{2}=$ $\left(\frac{r}{1-\left(r\eta\right)^{2}}\right)\left(-dt^{2}+dx^{2}\right)+r\left\{f_{s}(r)d\varphi^{2}+\frac{2}{e_{A}^{2}}\left[\hat{\omega}_{1}^{2}+\hat{\omega}_{2}^{2}+\left(\hat{\omega}_{3}-e_{A}Q_{A}\zeta(r)d\varphi\right)^{2}\right]\right$. (3.4)
@@
$B=$ $-\left(\frac{r}{1-\left(r\eta\right)^{2}}\eta r\right)dx\wedge dt$
$\Phi=$ $\frac{1}{2}\ln\left(\frac{r^{2}}{1-\left(\eta r\right)^{2}}\right).$
@@
The background generated is, as expected [15], [16], a supergravity solution of type IIB, as can be checked using the equations in the Appendix A.1. The field content in this supergravity setup implies the presence of D3-branes that couples to the $C_{4}$ field, D5-branes coupling magnetically to $C_{2}$ and F1-strings, coupling electrically with the Kalb-Ramond B field, with these coupling generating fluxes that are quantized, [24], [25]. For the D5-branes, they are the same as before deforming, $N_{A}=N_{A,D5}$ and $N_{B}=N_{B,D5}$. To obtain the Page charges, we follow the line of reasoning of [26], and with the calculations done in detail in Appendix A.3, we get
@@
We then choose the configuration containing strings at $\psi_{A}=e_{B}^{2}/\#$ and $\psi_{B}=e_{A}^{2}/\#$, where #\# stands for a normalization, since we know that we must have $\psi_{A,B}<4\pi$ (which we will ignore from now on). This specific configuration, as claimed before, is consistent with $\lim_{r\rightarrow\infty}\rho(r)=0$, where $\rho(r)$ is the radial effective density obtained after integrating the compact part of the cycle, $\rho(r)=\int d\Omega\rho_{1}$. In fact, we have
@@
As seen in [10, 1, 8], the dual gauge theory to the I and II backgrounds of [1] are supersymmetric and confining. Since the spacetime geometry corresponds to a BPS supergravity solution, the dual field theory is expected to be supersymmetric. Moreover, the linear dilaton—which diverges as $r\to\infty$—indicates a running coupling that increases with energy, a behavior typically associated with confinement in the dual gauge theory.
@@ 3.2 Penrose limit of the TsT transformation ($T\bar{T}$ deformation)
$\mathcal{L}=\frac{2r_{+}d\psi_{A}^{2}}{e_{A}^{2}}+\frac{2r_{+}d\psi_{B}^{2}}{e_{B}^{2}}-\frac{r_{+}dt^{2}}{1-\eta^{2}r_{+}^{2}}.$ (3.23)
@@
$\mathcal{L}=r_{+}\left(4d\psi^{2}-\frac{dt^{2}}{1-\eta^{2}r_{+}^{2}}\right).$ (3.24)
@@
$du$ $=2d\psi-\frac{dt}{\sqrt{1-\eta^{2}r_{+}^{2}}},$ (3.25)
$dv$ $=2d\psi+\frac{dt}{\sqrt{1-\eta^{2}r_{+}^{2}}}\;,$
@@
$L^{-2}ds^{2}={}$ $-\frac{r_{+}}{1-\eta^{2}r_{+}^{2}}dt^{2}+\frac{r_{+}}{1-\eta^{2}r_{+}^{2}}dx^{2}+d\varphi^{2}(r-r_{+})r_{+}f_{s}^{\prime}(r_{+})+\frac{4dr^{2}}{(r-r_{+})r_{+}f_{s}^{\prime}(r_{+})}$ (3.28)
@@
$x=\frac{X}{L},\;\psi_{A,B}=e_{A,B}\left(\frac{\Theta}{L}\pm\psi\right),\;t=\sqrt{1-\eta^{2}r_{+}^{2}}\left(\frac{v}{2}-\frac{u}{2L^{2}}\right)\;,$
@@
$+\frac{16d\rho^{2}}{L^{2}r_{+}f_{s}^{\prime}(r_{+})}+\frac{\rho^{2}r_{+}f_{s}^{\prime}(r_{+})d\varphi^{2}}{L^{2}}+\frac{r_{+}dX^{2}}{L^{2}(1-\eta^{2}r_{+}^{2})}$
@@
$\phi=\frac{1}{2}\log\frac{r_{+}^{2}}{1-\eta^{2}r_{+}^{2}}$ (3.33)
$B=\frac{\eta r_{+}^{2}}{2L\sqrt{1-\eta^{2}r_{+}^{2}}}dv\wedge dX$
@@
in the limit, $F_{(3)}$ is constant and finite, as we obtain correctly $\sim 1/L^{2}$, but $B\sim{\cal O}(1/L)\Rightarrow\infty$ (for finiteness, it should have been $\sim 1/L^{2}$). That seems like a problem, but that is a somewhat common occurrence (see for instance [5]): the point is that this $B$ field is an infinite constant, and as such is a pure gauge, that can be gauged away, so dropped. Note that the next term in the $1/L$ expansion would be ${\cal O}(1/L^{3})$, which vanishes in the $L\rightarrow\infty$ limit, so in the end we can put $B=0$.
@@
However, before we study the spectrum, we are going to do another change of coordinates. From (3.46), we can do the change of coordinates analogous to the one done in [4],
@@ 3.2.1 The operator-state mapping from string oscillators to the gauge theory
$p^{-}\equiv p^{u}=-p_{V}=\frac{\sqrt{1-\eta^{2}r_{+}^{2}}}{2}\Delta-\frac{J_{\psi}}{4}$ (3.64)
$p^{+}\equiv p^{V}=-p_{u}=-\frac{\frac{\sqrt{1-\eta^{2}r_{+}^{2}}}{2}\Delta+\frac{J_{\psi}}{4}}{L^{2}}\;,$
where $\Delta,J$ are eigenvalues of the energy $H$, corresponding to conformal dimension, and momentum $p_{\psi}=\frac{1}{2}(e_{A}p_{\psi_{A}}+e_{B}p_{\psi_{B}})$, corresponding to a $U(1)$ (global) R-charge, respectively. The negative sign here is a result of the choice in doing the change of coordinates in (3.25), and does not affect the results. Therefore, the Penrose limit corresponds to a specific sector of the gauge theory in which the surviving operators (à la BMN [3]) have large charge $J_{\psi}$, approximately equal to a large conformal dimension $\Delta$,
$J_{\psi}\simeq 2\sqrt{1-(\eta r_{+})^{2}}\Delta\propto L^{2}\;.$ (3.65)
@@
In 2+1 dimensions, a scalar has the classical (conformal) dimension 1/2, so we must define the unit of $J_{\psi}$ to be 1, so that $E=p^{-}=0$ for it. But, of course, the implicit assumption is that the TsT deformation by a parameter $\eta$ is such that $\Delta_{\eta}\sqrt{1-(\eta r_{+})^{2}}=1$, or $\Delta_{\eta}=1/\sqrt{1-(\eta r_{+})^{2}}$. We redefine the zero of the energy by an $E_{0}=1$, and then we find that, in order to get the correct spectrum, we must rescale $H$ by $\mu=-2$. Summarizing,
@@
Alternatively, more in line with what we obtained, we can define a unique vacuum (with no $a$ index), defined just by $Z^{J}$, and insertions of $W$ correspond to massless string oscillators $(x_{8}$) insertions. In the case of no string momentum, $n_{i}=0$, there is no difference between the two interpretations, but at $n_{i}\neq 0$, the latter interpretation should be the one that matches. However, we will not attempt here to match the $n_{i}\neq 0$ case, and we will leave it for further work. We just note here that the BMN operator analysis of the TsT deformed is case is essentially the same as in the undeformed case, with the deformation just acting on $\Delta_{\eta}$ and $J_{\psi}$ simultaneously, but otherwise keeping the solvability of the sector intact. The deformation should also modify the parameter $b$, which in gauge theory should be proportional to $g^{2}_{YM}\propto e^{\Phi}$, which we saw that gets modified by $1/\sqrt{1-\eta^{2}r_{+}^{2}}$, the same factor as in $\Delta_{\eta}$.

### v4

@@ Abstract
In this paper we analyze a generalized ”single-trace $T\bar{T}$” deformation, defined by a TsT transformation, of the fibered $I$-brane solution from [1]. We use the Penrose limit to try to understand it, and we consider both the TsT followed by the Penrose limit, as well as the Penrose limit followed by TsT. We describe the dual spin chains obtained in field theory. In the first case we find that, indeed, the TsT transformation preserves solvability in a simple way, as in the standard $T\bar{T}$ case. In the second case, we have several options, but none is simple enough to be conclusive, however, one case gives us an asymptotically free and IR nontrivial field theory sector, and another a new parallelizable pp wave.
@@ 2 Review of $I$-brane background(s)
The supersymmetry of this background was studied in [1], where it was found that it preserves four supersymmetries when the parameters $e_{A,B},m,Q_{A,B}$ satisfy $m=0,\ e_{A}Q_{B}=\pm e_{B}Q_{A}$. The background is confining, and the space ends smoothly (without a singularity) in the IR, at $r=r_{+}$, the largest of the solutions of $f_{s}(r)=0$. This is possible because of the compactification on the $S^{1}$ defined by $\varphi$. Since $e^{\Phi}$ will become large at some $r_{*}$, the above D5-D5 solution wrapped on $S^{3}$ and $\tilde{S}^{3}$ (as defined by the large $r$ region), needs to be S-dualized to an NS5-NS5 one, see [1, 10].
@@ 3.1 TsT transformation ($T\bar{T}$ deformation) in gravity and gauge theory
The general rules for T duality (both for the NSNS and the RR fields) are known from the literature, see [20], [21], [22], but for the reader’s convenience, the general rules are given in the Appendix A.2. However, as in [23, 24, 25], one needs to first Wick rotate to Euclidean space, then to the TsT transformation, and then Wick rotate back, in order not to get an apparent singularity in the metric. Since the TsT transformation acts on $x$ and $t$, in the intermediate steps we rewrite only the parts of the metric affected by these transformations. Performing first a T-duality along $x$, we obtain
$ds^{2}$ $=$ $+rdt^{2}+\frac{1}{r}dx^{2}+\cdots.$ (3.1)
@@
$ds^{2}=+rdt^{2}+2\eta rdtdx+\left(\frac{1+\left(r\eta\right)^{2}}{r}\right)dx^{2}+\cdots\;,$ (3.3)
@@
Finally we T-dualize back in $x$, and Wick rotate back to Minkowski space. The NSNS part of the background is
$ds_{st}^{2}=$ $\left(\frac{r}{1+\left(r\eta\right)^{2}}\right)\left(-dt^{2}+dx^{2}\right)+r\left\{f_{s}(r)d\varphi^{2}+\frac{2}{e_{A}^{2}}\left[\hat{\omega}_{1}^{2}+\hat{\omega}_{2}^{2}+\left(\hat{\omega}_{3}-e_{A}Q_{A}\zeta(r)d\varphi\right)^{2}\right]\right$. (3.4)
@@
$B=$ $+\left(\frac{r}{1+\left(r\eta\right)^{2}}\eta r\right)dx\wedge dt$
$\Phi=$ $\frac{1}{2}\ln\left(\frac{r^{2}}{1+\left(\eta r\right)^{2}}\right).$
Since the TsT transformation is a symmetry of the string theory, and it acts on the field theory coordinates, the modification should not change the solvability of the theory, a fact consistent with $T\bar{T}$. Also, since the deformation is ${\cal O}(r^{2})={\cal O}(r^{\Delta-d})$ (with $r$ the standard radial coordinate and $d=2$), this gives a $\Delta=4$ deformation, just like $T\bar{T}$ is supposed to be in $d=2$. Also, since it is the metric that is deformed, that couples to $T_{\mu\nu}$, it should be something like $\sim\det T_{\mu\nu}\sim T\bar{T}$. We will say more about this later, but this is similar to what was obtained in [25], just that now in the better understood $d=2$ case, for instance since [17] rigorously proved that TsT equals $T\bar{T}$ for the NS-NS $AdS_{3}\times S^{3}\times T^{4}$ case. Also, exactly like in that case, we note that, since $\Delta>d$, the deformation corresponds to an irrelevant operator, thus modifies the UV of the theory, which is not a CFT anymore (as it happens for any $T\bar{T}$ deformation).
@@
The background generated is, as expected [15], [16], a supergravity solution of type IIB, as can be checked using the equations in the Appendix A.1. The field content in this supergravity setup implies the presence of D3-branes that couples to the $C_{4}$ field, D5-branes coupling magnetically to $C_{2}$ and F1-strings, coupling electrically with the Kalb-Ramond B field, with these coupling generating fluxes that are quantized, [27], [28]. For the D5-branes, they are the same as before deforming, $N_{A}=N_{A,D5}$ and $N_{B}=N_{B,D5}$. To obtain the Page charges, [^2] To remind the reader, the Page charge of Dpp-branes, $Q_{\rm Page,p}=\frac{1}{(2\pi l_{s})^{7-p}}\int_{\Sigma_{8-p}}\hat{F}_{8-p}$, where $\hat{F}_{p}=[e^{-B\wedge}\tilde{F}]_{p}$, is the only one that is quantized, but it is not gauge invariant (changes under large gauge transformations), while the Maxwell charge, $Q_{\rm Max,p}=\frac{1}{(2\pi l_{s})^{7-p}}\int\tilde{F}_{8-p}$, is conserved and gauge invariant, but not quantized, where $\tilde{F}_{p}$ has a nontrivial Bianchi identity involving $H_{3}=dB$, for instance $d\tilde{F}_{4}=H_{3}\wedge\tilde{F}_{2}$, and the brane charge is one that is gauge invariant, but not conserved nor quantized, and is the integral of the source (breaking) of the Bianchi identity, for instance $\int_{M^{5}}(d\tilde{F}_{4}-H_{3}\wedge\tilde{F}_{2})$. For the F1, this is a bit less well defined, but we will follow [29] for the analog of the (quantized) Page charge. we follow the line of reasoning of [29], and with the calculations done in detail in Appendix A.3, we get
@@
We then choose the configuration containing strings at $\psi_{A}=e_{B}^{2}/\#$ and $\psi_{B}=e_{A}^{2}/\#$, where #\# stands for a normalization, since we know that we must have $\psi_{A,B}<4\pi$ (which we will ignore from now on, since we are not so much after an exact quantization condition, but after the existence or not of various branes in the geometry). This specific configuration, as claimed before, is consistent with $\lim_{r\rightarrow\infty}\rho(r)=0$, where $\rho(r)$ is the radial effective density obtained after integrating the compact part of the cycle, $\rho(r)=\int d\Omega\rho_{1}$. In fact, we have
@@
As seen in [10, 1, 8], the dual gauge theory to the I and II backgrounds of [1] are supersymmetric and confining. Since the spacetime geometry corresponds to a BPS supergravity solution, the dual field theory is expected to be supersymmetric. Moreover, the linear dilaton—which diverges as $r\to\infty$—indicates a running coupling that increases with energy, a behavior typically associated with asymptotic freedom in the dual gauge theory, while [1] showed that the smooth ending of the space in the IR (with cigar-like topology) does indeed lead to confinement, and that the gauge coupling, computed from the action of a probe 5-brane, behaves essentially like the dilaton. However, the asymptotic freedom is unlike the undeformed case of [1], where the coupling was bounded. See Appendix A.4 for details of the calculation.
@@ 3.2 Penrose limit of the TsT transformation ($T\bar{T}$ deformation)
$\mathcal{L}=\frac{2r_{+}d\psi_{A}^{2}}{e_{A}^{2}}+\frac{2r_{+}d\psi_{B}^{2}}{e_{B}^{2}}-\frac{r_{+}dt^{2}}{1+\eta^{2}r_{+}^{2}}.$ (3.23)
@@
$\mathcal{L}=r_{+}\left(4d\psi^{2}-\frac{dt^{2}}{1+\eta^{2}r_{+}^{2}}\right).$ (3.24)
@@
$du$ $=2d\psi-\frac{dt}{\sqrt{1+\eta^{2}r_{+}^{2}}},$ (3.25)
$dv$ $=2d\psi+\frac{dt}{\sqrt{1+\eta^{2}r_{+}^{2}}}\;,$
@@
$L^{-2}ds^{2}={}$ $-\frac{r_{+}}{1+\eta^{2}r_{+}^{2}}dt^{2}+\frac{r_{+}}{1+\eta^{2}r_{+}^{2}}dx^{2}+d\varphi^{2}(r-r_{+})r_{+}f_{s}^{\prime}(r_{+})+\frac{4dr^{2}}{(r-r_{+})r_{+}f_{s}^{\prime}(r_{+})}$ (3.28)
@@
$x=\frac{X}{L},\;\psi_{A,B}=e_{A,B}\left(\frac{\Theta}{L}\pm\psi\right),\;t=\sqrt{1+\eta^{2}r_{+}^{2}}\left(\frac{v}{2}-\frac{u}{2L^{2}}\right)\;,$
@@
$+\frac{16d\rho^{2}}{L^{2}r_{+}f_{s}^{\prime}(r_{+})}+\frac{\rho^{2}r_{+}f_{s}^{\prime}(r_{+})d\varphi^{2}}{L^{2}}+\frac{r_{+}dX^{2}}{L^{2}(1+\eta^{2}r_{+}^{2})}$
@@
$\phi=\frac{1}{2}\log\frac{r_{+}^{2}}{1+\eta^{2}r_{+}^{2}}$ (3.33)
$B=\frac{\eta r_{+}^{2}}{2L\sqrt{1+\eta^{2}r_{+}^{2}}}dv\wedge dX$
@@
in the limit, $F_{(3)}$ is constant and finite, as we obtain correctly $\sim 1/L^{2}$, but $B\sim{\cal O}(1/L)\Rightarrow\infty$ (for finiteness, it should have been $\sim 1/L^{2}$). That seems like a problem, but that is a somewhat common occurrence (see for instance [5]): the point is that this $B$ field is an infinite constant, and as such is a pure gauge, that can be gauged away, so dropped. Note that the next term in the $1/L$ expansion would be ${\cal O}(1/L^{3})$, which vanishes in the $L\rightarrow\infty$ limit, so in the end we can put B=0B=0.333Note that the equations of motion for the resulting solution have been checked.
@@
[…]
@@ 3.2.1 The operator-state mapping from string oscillators to the gauge theory
$p^{-}\equiv p^{u}=-p_{V}=\frac{\sqrt{1+\eta^{2}r_{+}^{2}}}{2}\Delta-\frac{J_{\psi}}{4}$ (3.64)
$p^{+}\equiv p^{V}=-p_{u}=-\frac{\frac{\sqrt{1+\eta^{2}r_{+}^{2}}}{2}\Delta+\frac{J_{\psi}}{4}}{L^{2}}\;,$
where $\Delta,J$ are eigenvalues of the energy $H$, corresponding to scaling dimension, and momentum $p_{\psi}=\frac{1}{2}(e_{A}p_{\psi_{A}}+e_{B}p_{\psi_{B}})$, corresponding to a $U(1)$ (global) R-charge, respectively. The negative sign here is a result of the choice in doing the change of coordinates in (3.25), and does not affect the results. Therefore, the Penrose limit corresponds to a specific sector of the gauge theory in which the surviving operators (à la BMN [3]) have large charge $J_{\psi}$, approximately equal to a large conformal dimension $\Delta$,
$J_{\psi}\simeq 2\sqrt{1+(\eta r_{+})^{2}}\Delta\propto L^{2}\;.$ (3.65)
@@
In 2+1 dimensions, a scalar has the classical (conformal) dimension 1/2, so we must define the unit of $J_{\psi}$ to be 1, so that $E=p^{-}=0$ for it. But, of course, the implicit assumption is that the TsT deformation by a parameter $\eta$ is such that $\Delta_{\eta}\sqrt{1+(\eta r_{+})^{2}}=1$, or $\Delta_{\eta}=1/\sqrt{1+(\eta r_{+})^{2}}$. We redefine the zero of the energy by an $E_{0}=1$, and then we find that, in order to get the correct spectrum, we must rescale $H$ by $\mu=-2$ (this is the same procedure as was needed in [6, 7, 8]). Summarizing,
@@
Alternatively, more in line with what we obtained, we can define a unique vacuum (with no $a$ index), defined just by $Z^{J}$, and insertions of $W$ correspond to massless string oscillators $(x_{8}$) insertions. In the case of no string momentum, $n_{i}=0$, there is no difference between the two interpretations, but at $n_{i}\neq 0$, the latter interpretation should be the one that matches. However, we will not attempt here to match the $n_{i}\neq 0$ case, and we will leave it for further work, as it is complicated, but will not change the $n_{i}=0$ matching for the (second) unique vacuum identification, as in, for instance, [4, 5] (also [6, 8, 7]). We just note here that the BMN operator analysis of the TsT deformed is case is essentially the same as in the undeformed case, with the deformation just acting on $\Delta_{\eta}$ and $J_{\psi}$ simultaneously, but otherwise keeping the solvability of the sector intact. The deformation should also modify the parameter $b$, which in gauge theory should be proportional to $g^{2}_{YM}\propto e^{\Phi}$, which we saw that gets modified by $1/\sqrt{1+\eta^{2}r_{+}^{2}}$, the same factor as in $\Delta_{\eta}$. What we could learn in this case is exactly what one was able to learn in [6, 8, 7] (and in [4, 5]).